OpenStudy (anonymous):

What's it mean when it says graph y=2x-4?

OpenStudy (anonymous):

it means graph the line using y=mx+b, where m=to the slope of the line, and b = to the y-intercept.

OpenStudy (anonymous):

yes now im just confused.

OpenStudy (anonymous):

okay, so are you familiar with the behavior of lines?

OpenStudy (anonymous):

get your paper with a coordinate graph, take your equation, y=2x-4. let x=0, so y=2(0)-4,and you will find that when x=0, y=-4. So you have a point (0,-4). Now let y=0, 0=2x-4,and you will find when y=0, x=2. So now this second point is (2,0). If you plot those points, and then connect the dots, you will have a line which has a slope that rises two units up, and to the right over one. So point (1,-2) would be in between those two point you had before.